# MegaScale-Infer: Serving Mixture-of-Experts at Scale with Disaggregated Expert Parallelism

### v3

@@ 1 Introduction
As the effectiveness of LLMs increasingly depends on the escalation of model parameters, there is a growing imperative to scale up these models [42, 48]. Due to the sparse activation architecture, mixture-of-experts (MoE) models [45, 55] are a practical choice for scaling. MoE dynamically route input tokens to a subset of feed-forward networks (FFNs), which are known as experts, rather than engaging all FFNs (i.e., all parameters). This design enables sub-linear scaling of required FLOPs as the number of experts and model size increases, significantly reducing computational complexity without compromising model quality.
Unfortunately, reduced computational complexity does not necessarily translate into lower computational costs in practical serving scenarios. This discrepancy arises from the mismatch between the characteristics of LLM inference and the compute capabilities of GPUs, a problem that becomes increasingly pronounced with growing MoE sparsity. Figure 1 demonstrates this issue. Specifically, an LLM consists of multiple layers of attention and FFN modules. During the decoding phase, which dominates the LLM inference process [44], the GPU utilization of attention modules remains low because they must access the intermediate states (i.e., key-value cache) of all previous tokens. Conversely, FFN modules achieve high GPU utilization as the number of tokens involved in computation increases.
@@
A natural solution is to disaggregate attention from the LLM inference process and replicate attention modules to increase the decoding batch size for FFN modules. This approach is adopted by Infinite-LLM [47], which focuses on optimizing dense model inference in long-context scenarios. In such cases, GPU memory capacity, rather than sparsity, is the primary constraint, and the communication pattern is relatively simple compared to the top-$k$ selection in MoE. Consequently, its solution is less effective in addressing the unique challenges of MoE inference.
We present MegaScale-Infer, an efficient and cost-effective system designed for large-scale MoE serving. MegaScale-Infer disaggregates the attention and expert modules, assigning them to separate GPUs—a strategy we term disaggregated expert parallelism. Our approach offers two major benefits. First, it enables independent scaling of each module with customized model parallelism strategies. Specifically, attention modules are replicated using data parallelism, while FFN modules are scaled with expert parallelism. By consolidating requests from multiple attention replicas, the GPU utilization of each expert increases significantly as the batch size per attention replica grows. Second, it enables the deployment of attention and FFN modules on heterogeneous GPUs to fully leverage their different capabilities and achieve lower costs. For example, attention modules can be deployed on GPUs with more cost-effective memory capacity and bandwidth, while FFN modules can utilize GPUs with more affordable compute capability. As shown in Figure 1(c), FFN can easily become compute-intensive in MegaScale-Infer, while attention achieves higher GPU utilization per cost under heterogeneous deployment.
@@
We implement MegaScale-Infer and evaluate it using MoE models with sizes ranging from 132 to 317 billion parameters. The experimental results show that MegaScale-Infer outperforms state-of-the-art LLM serving systems by up to 1.9$\times$ in per-GPU decoding throughput. We also conduct experiments on a heterogeneous cluster, where MegaScale-Infer achieves 1.7$\times$ higher throughput per dollar. Compared to NCCL [6], a widely-used communication library, MegaScale-Infer’s M2N communication achieves 4.2$\times$ higher throughput and 68.2% lower latency.
@@
  • Our experiments demonstrate significant improvements in throughput and cost-effectiveness when utilizing our system’s unique capabilities. MegaScale-Infer achieves up to 1.90$\times$ and 1.86$\times$ per-cost decoding throughput against state-of-the-art LLM serving systems on homogeneous and heterogeneous clusters, respectively.
@@ 2.1 LLM Inference Characteristics
The autoregressive generation pattern makes the attention module compute-intensive during the prefill phase and memory-intensive during the decoding phase. Even with request batching [72, 16], a widely-used optimization in efficient LLM serving, attention during the decoding phase remains the same memory access intensity. This is because each request has its own KV cache of input and previously generated tokens, which is different from each other. In the decoding iteration, each request must access its respective KV cache. In contrast, the FFN module’s computation only requires loading the corresponding model weights from GPU memory to SRAM, which can be shared across all tokens from different requests. Consequently, as presented in Figure 1(a), batching is only efficient for FFNs to reuse model parameters and improve GPU utilization.
@@ 2.2 LLM Serving at Scale
The scaling law [43] highlights the significance of model size as a key determinant of the model capability. To achieve state-of-the-art model capability, many efforts [42, 48] have been invested in scaling LLMs up to hundreds of billions of parameters. Due to the large model size, serving these models necessitates both algorithmic and system optimizations.
@@ 2.3 Problems in Large-scale MoE Serving
Take Mixtral 8x22B as a more concrete example. Assume that we use NVIDIA A100-SXM-80GB GPUs, which have a computational power of 312 TFLOPS and memory bandwidth of 2 TB/s, to serve this model with the bfloat16 datatype. The floating point operations required for a $b\times h$ to $h\times n$ GEMM (General Matrix to Matrix Multiplication) are $2bhn$, where $b$ and $h$ represent the decoding batch size and the model’s hidden dimension size, respectively. The number of parameters this GEMM needs to access is $hn$, and the data volume is $2hn$ for bfloat16. Let the GPU’s floating point compute capability be $F$ and the memory bandwidth be $B$. According to the roofline model [67], a GPU requires that $\frac{2bhn}{F}\geq\frac{2hn}{B}$, i.e., $b\geq\frac{F}{B}$, to fully utilize its matrix multiplication capability. For an A100 GPU, the batch size at least needs to be 156 tokens ($\frac{312TFLOPS}{2TB/s}$). However, given a batch size of 156, the average number of decoding tokens dispatched to each expert is $39156\times topk/\#expert=156\times 2/8=$ 39, with the theoretical Model Flops Utilization (MFU) for FFN modules of $2/8=25\%$. Formally, the theoretical relationship between batch size and FFN’s GPU utilization for dense models is $util=min(\frac{B}{F}b,1)$, but for MoE, it is $util=min(\frac{topk}{\#expert}\frac{B}{F}b,1)$.
@@ 2.4 Opportunities and Challenges
  • Heterogeneous deployment. The disaggregated architecture naturally separates the deployment for attention and FFN modules, allowing for the use of most cost-effective GPUs for each. It also opens up opportunities to use specialized hardware and software to separately accelerate attention and FFN computation.
@@ 3 MegaScale-Infer Overview
Figure 3: MegaScale-Infer runtime instance architecture.
@@
$h^{\prime}$ Intermediete dimension size of FFN
@@
$tpd$ throughput per dollar
@@
High-performance M2N communication. MegaScale-Infer employs a customized M2N communication library to transfer the intermediate outputs between each pair of attention nodes and expert nodes. To achieve efficient data transmission, the library removes unnecessary GPU-to-CPU data copies, group initialization overhead, and GPU synchronization. It also proposes traffic-oriented optimizations specific to this scenario.
@@ 4 Disaggregated Expert Parallelism
In this section, we present the design of ping-pong pipeline parallelism and the approach to generating the deployment plan of MegaScale-Infer. Given the MoE model, workload characteristics (e.g., sequence lengths), available hardware, and latency requirements, MegaScale-Infer determines the deployment plan by specifying $(i)$ the respective parallelism strategies for attention and experts, $(ii)$ the number of micro-batches for the ping-pong pipeline, $(iii)$ the maximum batch size, and $(iv)$ the hardware setup for deployment. Our goal is to identify the deployment plan that maximizes throughput per unit cost (e.g., dollar). Table 1 lists the key notations in our discussion. We assume the model uses grouped-query attention (GQA) [19], which is the most popular method for attention.
@@ 4.1 Ping-Pong Pipeline Parallelism
Constraint 1 aims to minimize the GPU idle time caused by computation dependencies across MoE layers. Constraint 2 and constraint 3 describe methods for hiding communication overhead. Specifically, the communication time for a single micro-batch must be shorter than the forward computation time of attention and experts, and the forward time of one MoE layer for the global batch on each node must be sufficient to cover the time required for a single micro-batch to pass through the layer. We can then obtain the minimum number of micro-batches needed using formula $m\geq 2\times(1+\frac{T_{c}}{T_{f}})$, where $10<\frac{T_{c}}{T_{f}}<$ 1. For deployments with fast communication ($T_{c}<\frac{1}{2}T_{f}$), at least 3 micro-batches are required. For those with relatively slower communication, at least 4 micro-batches are required.
@@ 4.2 Deployment Plan Search
8:                  $B$, $tpd\leftarrow$ simulate($G,plan,SLO$)
9:                  $plan\leftarrow plan\cup\{B,tpd\}$
10:                  if $plan^{*}.tpd<plan.tpd$ then
@@
Considering ping-pong pipeline parallelism, the search space of MegaScale-Infer deployment plan includes the tensor parallelism sizes for attention nodes ($tp_{a}$) and expert nodes ($tp_{e}$), the number of attention nodes ($n_{a}$), the number of micro-batches, and the global batch size ($B$). Our objective is to minimize the throughput per unit cost while adhering to the SLO constraint. Algorithm 1 shows the pseudo-code for searching the optimal deployment plan given hardware setup and model configurations. It enumerates the feasible $tp_{a}$ and $tp_{e}$, subject to GPU memory capacity limit. For each pair of $tp_{a}$ and $tp_{e}$, it calculates the number of attention nodes to balance the computation time as closely as possible according to constraint 1. The algorithm then compares the throughput per unit cost among deployment plans with varying numbers of micro-batches. Using the SIMULATE function, it finds the maximum global batch size that meets the SLO through binary search and obtains the optimal plan.
The complexity of Algorithm 1 is $O(M^{2}N_{m})$, with $M$ as the GPU limit per server and $N_{m}$ as the maximum number of micro-batches. Typically, $M$ has four choices (e.g., $\{1,2,4,8\}$) in modern clusters. We set $N_{m}$ to four because splitting into too many micro-batches reduces GEMM efficiency in expert nodes and thus increases the latency. Consequently, the search space remains manageable.
@@
Constraint 8 represents the GPU memory capacity limit for (bfloat16) KV cache size. And the throughput per dollar is $\frac{B/T_{total}}{tp_{a}n_{a}Cost_{a}+tp_{e}ECost_{e}}$.
4.3 Heterogeneous Deployment
Table 3 lists the performance specifications, prices, and corresponding ratios for a selection of NVIDIA GPUs. We enumerate the scenarios of using each type of GPU as the hardware for attention or expert nodes to determine the optimal deployment plan. Intuitively, H20 is more suitable for attention due to its large memory capacity and high memory bandwidth per cost. Meanwhile, the L40S GPU is more cost-effective for experts.
Accelerator Price Cap. (GB) Bw. (GB/s) Comp. (TFLOPS) Performance per Cost
@@ 5 High-Performance M2N Communication
  • Instability at higher percentiles. The performance issue highlighted in Figure 5(a) consistently exacerbates at higher percentiles, as shown in Figure 5(b). At the 99th percentile, the baseline experiences only a slight increase in latency, whereas NCCL exhibits a significant surge, particularly when scaling to 32 receivers.
The underlying causes of these issues are multifaceted, stemming from specific design choices in NCCL that are not well-suited for this particular use case. Regarding overhead, we first note that NCCL networking requires intermediate copies [9] to transfer data from the GPU memory to the CPU proxy, which performs the network operations. While features like user buffer registration [7] aim to reduce these copies, they do not fully eliminate them. Second, peer-to-peer group operations [8] are processed in batches of at most $8$ operations, which causes a harmful effect as the number of receivers scales. Third, as a general-purpose collective communication library, NCCL incurs overhead from general group operation setup, including preparing and launching a batch of N send operations, internal handling and verifications, etc. While these steps are essential for ensuring broad applicability, they introduce unnecessary latency and can be optimized, though not entirely eliminated. Regarding stability, this issue is much more complex and can arise from multiple sources, including OS, memory, networking, and GPU thermal differences [59, 71, 21, 36, 32, 62, 73]. Previous studies have highlighted that common sources of instability often arise from GPU synchronization operations and device memory accesses [75, 34], both of which are prevalent in NCCL but absent in the baseline.
@@
M2N Sender. Figure 6 depicts the components of an M2N sender. In order to comply with the stream-oriented programming model, ① M2N senders launch low-utilization kernels which ensures that the ② pre-registered tensor to be transmitted is properly populated by preceding computation kernels. ③ Once the tensor is ready, ④ our in-house CPU communication library (denoted as Core Sender in the figure) transmits the tensor efficiently using RDMA write with immediate [11]. Finally, to guarantee proper utilization of the registered tensor, ⑤ the sender polls completions from the corresponding completion queue  [10], confirming that data has been written to the remote buffers. This design eliminates complex GPU synchronization, GPU-to-CPU copies, and group initialization overhead, all of which contribute to significant latency issues, especially for relatively small tensor sizes, as demonstrated in Figures 10 and 11.
M2N Receiver. Figure 7 illustrates the components of an M2N Receiver. Since communication between M2N senders and receivers leverages ⑥ pre-registered RDMA buffers, the receiver’s primary tasks are relatively straightforward. First, it must verify that tensors have been successfully transmitted from the corresponding senders, which is efficiently achieved by ⑦ polling the relevant completion queue. Second, to ensure data consistency at the GPU level, our in-house CPU communication library (denoted as Core Receiver in the figure) leverages GDRCopy [12] and performs a ⑧ flush operation [1]. Finally, the receiver ensures that the subsequent kernel in the compute stream waits until the data has been correctly placed in the destination buffer, which is achieved by ⑨ updating a shared flag. This simpler design eliminates the need for GPU-to-GPU copies and effectively reduces the GPU utilization overhead of receivers.
@@
  • Congestion control fine-tuning. We observed substantial latency degradations in unbalanced communication scenarios, where the amount of data to be sent varies significantly per receiver. To address this, we fine-tune our congestion control algorithms to minimize rate-limiting effects and allow faster convergence in general.

### v4

@@ 1 Introduction
As the effectiveness of LLMs increasingly depends on the escalation of model parameters, there is a growing imperative to scale up these models [49, 35]. Due to the sparse activation architecture, mixture-of-experts (MoE) models [52, 62] are a practical choice for scaling. MoE dynamically routes input tokens to a subset of feed-forward networks (FFNs), which are known as experts, rather than engaging all FFNs (i.e., all parameters). This design enables sub-linear scaling of required FLOPs as the number of experts and model size increases, significantly reducing computational complexity without compromising model quality.
Unfortunately, reduced computational complexity does not necessarily translate into lower computational costs in practical serving scenarios. This discrepancy arises from the mismatch between the characteristics of LLM inference and the compute capabilities of GPUs, a problem that becomes increasingly pronounced with growing MoE sparsity. Figure 1 demonstrates this issue. Specifically, an LLM consists of multiple layers of attention and FFN modules. During the decoding phase, which dominates the LLM inference process [51], the GPU utilization of attention modules remains low because they must access the intermediate states (i.e., key-value cache) of all previous tokens. Conversely, FFN modules achieve high GPU utilization as the number of tokens increases.
@@
[…]
We present MegaScale-Infer, an efficient and cost-effective system designed for large-scale MoE serving. MegaScale-Infer disaggregates the attention and expert modules, assigning them to separate GPUs—a strategy we term disaggregated expert parallelism. Our approach offers two major benefits. First, it enables independent scaling of each module with customized model parallelism strategies. Specifically, attention modules are replicated using data parallelism, while FFN modules are scaled with expert parallelism. By consolidating requests from multiple attention replicas, the GPU utilization of each expert increases significantly as the batch size per attention replica grows. Second, it enables the deployment of attention and FFN modules on heterogeneous GPUs to fully leverage their different capabilities and achieve lower costs. For example, attention modules can be deployed on GPUs with more cost-effective memory capacity and bandwidth, while FFN modules can utilize GPUs with more affordable compute capability. As shown in Figure 1(c), FFN can easily become compute-intensive in MegaScale-Infer, while attention achieves higher GPU utilization per unit cost under heterogeneous deployment.
@@
We implement MegaScale-Infer and evaluate it using MoE models with sizes ranging from 132 to 317 billion parameters. The experimental results show that MegaScale-Infer outperforms state-of-the-art LLM serving systems by up to 1.9$\times$ in per-GPU decoding throughput. We also conduct experiments on a heterogeneous cluster, where MegaScale-Infer achieves 1.7$\times$ higher throughput per unit cost. Compared to NCCL [8], a widely-used communication library, MegaScale-Infer’s M2N communication achieves 4.2$\times$ higher throughput and 68.2% lower latency. MegaScale-Infer has already been deployed in the company’s inference services and reduces the serving cost by 1.5–2.0$\times$.
@@
  • Our experiments demonstrate significant improvements in throughput and cost-effectiveness with our system’s unique capabilities. MegaScale-Infer achieves up to 1.90$\times$ and 1.86$\times$ per-cost decoding throughput against state-of-the-art LLM serving systems on homogeneous and heterogeneous clusters, respectively.
This work does not raise any ethical issues.
@@ 2.1 LLM Inference Characteristics
The autoregressive generation pattern makes the attention module compute-intensive during the prefill phase and memory-intensive during the decoding phase. Even with request batching [79, 22], a widely-used optimization in efficient LLM serving, attention during the decoding phase remains the same memory access intensity. This is because each request has its own KV cache of input and previously generated tokens, which is different from each other. In the decoding iteration, each request must access its respective KV cache. In contrast, the computation of FFN only requires loading the corresponding model weights from GPU memory to SRAM, which can be shared across all tokens from different requests. Consequently, as presented in Figure 1(a), batching is only efficient for FFNs to reuse model parameters and improve GPU utilization.
@@ 2.2 LLM Serving at Scale
The scaling law [50] highlights the significance of model size as a key determinant of the model capability. To achieve state-of-the-art model capability, many efforts [49, 35] have been invested in scaling LLMs to hundreds of billions of parameters. Due to the large model size, serving these models necessitates both algorithmic and system optimizations.
@@ 2.3 Problems in Large-scale MoE Serving
[…]
@@
Figure 3: MegaScale-Infer runtime instance architecture.
@@ 2.4 Opportunities and Challenges
  • Heterogeneous deployment. The disaggregated architecture naturally separates the deployment for attention and FFN modules, allowing for the use of the most cost-effective GPUs for each. It also opens up opportunities to use specialized hardware and software to separately accelerate attention and FFN computation.
@@ 3 MegaScale-Infer Overview
$h^{\prime}$ Intermediate dimension size of FFN
@@
$tpuc$ throughput per unit cost
@@
High-performance M2N communication. MegaScale-Infer employs a customized M2N communication library to transfer the intermediate outputs between each pair of attention nodes and expert nodes. To achieve efficient and stable data transmission, the library removes unnecessary GPU-to-CPU data copies, group initialization overhead, and GPU synchronization. It also proposes traffic-oriented optimizations specific to this scenario.
@@ 4 Disaggregated Expert Parallelism
[…]
@@ 4.1 Ping-Pong Pipeline Parallelism
Constraint 1 aims to minimize the GPU idle time caused by computation dependencies across MoE layers. Constraint 2 and constraint 3 describe methods for hiding communication overhead. Specifically, the communication time for a single micro-batch must be shorter than the forward computation time of attention and experts, and the forward time of one MoE layer for the global batch on each node must be sufficient to cover the time required for a single micro-batch to pass through the layer. We can then obtain the minimum number of micro-batches needed using formula $m\geq 2\times(1+\frac{T_{c}}{T_{f}})$, where $0<\frac{T_{c}}{T_{f}}<1$. For deployments with fast communication ($T_{c}<\frac{1}{2}T_{f}$), at least 3 micro-batches are required. For those with relatively slower communication, at least 4 micro-batches are required.
@@
[…]
9:      $plan\leftarrow plan\cup\{B,tpuc\}$
10:      if $plan^{*}.tpuc<plan.tpuc$ then
@@
4.2 Deployment Plan Search
Considering ping-pong pipeline parallelism, the search space of MegaScale-Infer deployment plan includes the tensor parallelism sizes for attention nodes ($tp_{a}$) and expert nodes ($tp_{e}$), the number of attention nodes ($n_{a}$), the number of micro-batches, and the global batch size ($B$). Our objective is to minimize the throughput per unit cost while adhering to the SLO constraint. Algorithm 1 shows the pseudo-code for searching the optimal deployment plan given hardware setup and model configurations. It enumerates the feasible $tp_{a}$ and $tp_{e}$, subject to GPU memory capacity limit. For each pair of $tp_{a}$ and $tp_{e}$, it calculates the number of attention nodes to balance the computation time as closely as possible according to constraint 1. The algorithm then compares the throughput per unit cost among deployment plans with varying numbers of micro-batches. Using the SIMULATE function, it determines the maximum global batch size that meets the SLO through binary search and obtains the optimal plan.
[…]
@@ 4.2 Deployment Plan Search
Constraint 8 represents the GPU memory capacity limit for bfloat16 KV cache size. And the throughput per unit cost is $\frac{B/T_{total}}{tp_{a}n_{a}Cost_{a}+tp_{e}ECost_{e}}$.
Accelerator
Price
Cap.
Bw.
(GB/s)
Comp.
Performance per Cost
@@
4.3 Heterogeneous Deployment
Table 3 lists the performance specifications, prices, and corresponding ratios for a selection of NVIDIA GPUs. We enumerate the scenarios of using each type of GPU as the hardware for attention or expert nodes to determine the optimal deployment plan. Intuitively, H20 is more suitable for attention due to its large memory capacity and high memory bandwidth per unit cost. Meanwhile, the L40S GPU is more cost-effective for experts.
Heterogeneous deployment can also reduce energy consumption by utilizing hardware with lower energy consumption per unit of compute or bandwidth. For example, the H20 and L40S GPUs have maximum power consumptions of 500W and 350W, respectively. Under the same power consumption, the L40S offers higher compute performance, while the H20 provides higher bandwidth. We demonstrate the improvements in both cost and energy efficiency achieved through heterogeneous deployment in §7.2.
@@ 5 High-Performance M2N Communication
  • Instability at higher percentiles. The performance issue highlighted in Figure 5(a) consistently exacerbates at higher percentiles, as shown in Figure 5(b). At the 99th percentile, the baseline experiences only a slight increase in latency, whereas NCCL exhibits a significant surge, particularly when scaling to 32 or more receivers.
The underlying causes of these issues are multifaceted, stemming from specific design choices in NCCL that are not well-suited for this particular use case. Regarding overhead, we first note that NCCL networking requires intermediate copies [11] to transfer data from the GPU memory to the CPU proxy, which performs the network operations. While features like user buffer registration [9] aim to reduce these copies, they do not fully eliminate them. Second, peer-to-peer group operations [10] are processed in batches of at most 888 operations, which causes a harmful effect as the number of receivers scales. Third, as a general-purpose collective communication library, NCCL incurs overhead from general group operation setup, including preparing and launching a batch of N send operations, internal handling and verifications, etc. While these steps are essential for ensuring broad applicability, they introduce unnecessary latency and can be optimized, though not entirely eliminated. Regarding stability, this issue is much more complex and can arise from multiple sources, including OS, memory, networking, and GPU thermal differences [66, 78, 26, 43, 39, 69, 81]. Previous studies have highlighted that common sources of instability often arise from GPU synchronization operations and device memory accesses [83, 41], both of which are prevalent in NCCL but absent in the baseline.
@@
M2N Sender. Figure 6 depicts the components of an M2N sender. In order to comply with the stream-oriented programming model, ① M2N senders utilize CUDA events [12] to wait for previous kernels and make sure that the ② pre-registered tensor to be transmitted is properly populated. Then, to ensure the next kernel in the stream starts after the transmission is complete, M2N senders utilize CUDA driver operations [15] to ③ block the stream. Once the stream is blocked, the data transmission proceeds in two steps: First, ④ our in-house CPU communication library (denoted as Core Sender in the figure) transmits the tensor efficiently using RDMA write with immediate [14]. Second, to guarantee proper utilization of the registered tensor, ⑤ the sender polls completions from the corresponding completion queue  [13], confirming that data has been written to the remote buffers. Finally, M2N senders ⑥ unblock the stream by updating a shared memory flag, allowing other kernels to continue reusing the registered memory. This design eliminates complex GPU synchronization, GPU-to-CPU copies, and group initialization overhead, all of which contribute to significant latency issues, especially for relatively small tensor sizes, as demonstrated in Figures 11 and 12.
@@
M2N Receiver. Figure 7 illustrates the components of an M2N Receiver. Just like its peer component, M2N receivers also need to adhere to the stream-oriented programming model. Specifically, they ① wait on CUDA events to ensure that the ② pre-registered tensor is no longer in use, and ③ block the stream to guarantee that subsequent kernels do not proceed until the operation is complete. Once the stream is blocked, the data collection proceeds in two steps: First, receivers must verify that data has been successfully transmitted from the corresponding senders, which is efficiently achieved by ④ polling the relevant completion queue. Second, to ensure data consistency at the GPU level, our in-house CPU communication library (denoted as Core Receiver in the figure) leverages GDRCopy [16] and performs a ⑤ flush operation [1]. Finally, M2N receivers ⑥ unblock the stream by updating a shared memory flag, allowing other kernels to continue utilizing the registered memory. This simpler design eliminates the need for GPU-to-GPU copies and effectively reduces the GPU utilization overhead of receivers.
@@
  • Congestion control fine-tuning. We observed substantial latency degradations in unbalanced communication scenarios, where the amount of data to be sent varies significantly per receiver. To address this, we fine-tune our congestion control algorithms to minimize rate-limiting effects and allow faster convergence.
Comparison with DeepEP. DeepEP [35, 5, 80] proposed by DeepSeek also optimizes network communication for large-scale expert parallel serving in MoE. The key difference between our approach and DeepEP lies in the communication strategy: we leverage CPUs for inter-node communication, whereas DeepEP employs direct GPU-to-GPU communication without the involvement of CPU proxies [2]. GPU-to-GPU communication approaches consume GPU compute resources on both the sender and receiver sides [19]. In contrast, our CPU-to-CPU communication avoids the need for GPU compute resource allocation, thereby allowing compute kernels to fully utilize the GPU. Moreover, GPU-to-GPU communication demands careful orchestration to mitigate contention between communication and computation kernels. DeepEP employs custom PTX (assembly-like) instructions [18] to minimize L2 cache usage—a resource shared between kernel types. Our design, by comparison, does not require such low-level optimizations. When handling requests within a single QP, the CPU achieves lower latency in our approach because its higher clock speed allows it to issue doorbells faster than the GPU. However, the GPU offers stronger parallel processing capabilities, with multiple SMs able to independently manage QPs. Consequently, DeepEP’s approach can achieve a higher packet transmission rate at the cost of GPU SM resources and may yield better throughput when the packet size is very small. In our scenario, the amount of data transferred between each sender-receiver pair typically reaches several hundred kilobytes (§7.3). At this scale, a single-threaded CPU is sufficient to saturate the bandwidth. If the number of experts increases further and the per-connection communication volume becomes smaller, leveraging the GPU’s superior parallel processing capabilities may offer greater advantages in terms of throughput.